OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Eight divided by x squared. + 4

OpenStudy (anonymous):

y=(8/x^2)+4??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[y = \frac{ 8 }{ x ^{2} } + 4\]

OpenStudy (anonymous):

f(x)=4 x nd g(x)=(2/x^2 )+ 1

OpenStudy (anonymous):

dont you want the f(x) to be *2/x^2) + 1 ?

OpenStudy (anonymous):

jk